# DiLi: A Lock-Free Asynchronously Distributable Linked List

### v1

@@ 62           Lines 2 to 2;
257
Algorithm 3 Operations: Insert and Split
5.3 The Split Algorithm
The Split operation (Algorithm˜3) takes two inputs – a sublist entry and an item (referred to as sItem) at which the sublist is to be split. This truncates the original sublist to a smaller key range and creates a new sublist for the remaining key range. The Split is performed in through 4 major steps, while always guaranteeing that the client operations stay lock-free:
  1. (1)
    Allocate new counters stCt and endCt (lines 3 and 3) for the new sublist. The purpose of these counts is to ensure that if a client operation overlaps with Split/Move, it is detected so that Split/Move can take corrective action.
  2. (2)
    Insert a connected block of a subtail followed by a subhead (as in Lines 3 and 3) at the next pointer of sItem (insertion is in Line 3). If sItem gets deleted before this, split operation fails (Line 3) and can be retried with another call at a different sItem.
  3. (3)
    Update the stCt and endCt pointers of the second half of the sublist (Lines 3 to 3). Note that if an item gets inserted and gets missed in the traversal, it must have copied the counters from the item it was previous to, whose counters (or their ancestors’) were updated during this traversal. Thus, they must also contain the new counter pointers.
  4. (4)
    Update the registry by adding an entry for the second half of the sublist, and then truncating the first half to its reduced keyMax (Lines 3 to 3).
The above approach has two corner cases to handle due to the asynchronous insertion. First, for client operations that concurrently search for a key higher than sItem, the search can get truncated at the inserted subtail in case the client operation traversal encounters it. For this purpose, the subtails contain a keyMax element stored. Specifically, the red lines from 2 to 2 check whether the key belongs to the first half of the sublist that is undergoing a concurrent split or to the second half. If it is in the second half, then its subhead pointer is updated, speeding up potential retries of the lock-free Harris List traversal.
Secondly, we can no longer expect the difference between stCt and endCt of the individual sublists to be 0 when there are no ongoing updates (insert/remove). This is because, during the counter updates to the second half of the sublist, some updates could have incremented the stCt of the left half and updated the endCt of the right half. This accumulated change to the counter difference when the sublist is static is denoted as the offset of the sublist. As would be explained later, the new offset is important for Move and Switch operations. To compute this new offset, we employ a spin (Lines 3 to 3) that breaks only when the computed values $a_{1}$ and $a_{2}$ reflect the new right half offset and left half offsets respectively. The condition works because the difference between the stCt and endCt of any sublist stabilizes to a minimum only when there are no ongoing updates in the sublist. This is true for both the left and right halves of the sublist, and since the increments went to exactly one of the two stCt counters and one of the two endCt counters for each update, the sum of the new offsets must be equal to the old offset when the two new sublists together have no ongoing updates. We note that this spin does not delay the client operations, and only delays background operations on the split sublists.
5.4 Move, Replay and Switch Algorithms
A Move routine invoked on a server machine (henceforth $S_{1}$) takes the registry entry of a sublist and the machine to move it to (henceforth $S_{2}$), as parameters. The goal of Move is to create a clone of a sublist and then make it the live copy. To achieve this, the Move routine as described in Algorithm˜4 traverses each item (by recursively calling the MoveItem routine) in the sublist in $S_{1}$ and creates a copy of them in $S_{2}$. The reference in $S_{2}$ of last moved item is sent back each time so that subsequent items can be linked to the previously moved item.
Move compensates for concurrent updates that the Move routine did not witness during its traversal, by 1. detecting that such concurrent updates occurred, and 2. replicatingthose updates correctly on $S_{2}$. The former is achieved by using stCt (incremented before the start of update operation) and endCt (incremented after the update operation); and the latter is achieved by sending replicates (Line 4 in Algorithm˜2 and Line 3 in Algorithm˜3) of these operations to $S_{2}$. As discussed in Section˜4, the inserts and removes will need further specification for the correct replay of the updates in $S_{2}$, as they may appear out of order when received. To account for this, the replay models the changes to the sublist as a result of the following update operations: InsertAfter(Ref, Item) and Delete(Ref), that distinctly convey where the insert CAS (Line 3) and remove CAS (Line 2) that had succeeded. If multiple items got inserted at a particular Ref $A$, that had been moved by the Move routine, then the item with the lower timestamp got inserted first. As for the Delete, the same item that was moved to $S_{2}$ is the one that gets deleted, and a subsequent insertion for the same key is treated as a different item. Since Ref present in $S_{1}$ for an item is not known by Ref copy of it in $S_{2}$, we denote the uniqueness of an item by the <sId, timestamp> tuple.
Using the above observations, we developed our Replay algorithm presented in Lines 4 to 4. The received move of an item and a received insert replicate utilize this routine. Once the prev item is identified using their <<sId, timestamp>> tuple, the Replay algorithm takes this prev item and traverses from prev to find the right position to insert the item. To not miss any item, the Move routine even moves marked items as part of the algorithm, which are then delinked once the cloned sublist becomes active. Because replay is concurrently invoked by Move and multiple Replicate threads, insertions at replay are retried. An extensive proof of correctness for the replay is provided in Section˜D.3.
Finally, to detect that a sublist is live in $S_{1}$, we utilize the stCt and endCt values of the sublist. All updates increment the stCt before performing the update, and increment the endCt after performing the update. However, if a replicate message is sent, the endCt is incremented only after the completion of the replay (Line 4 and 4). This way, when the endCt measured matches the stCt measured immediately after, we can be sure that both copies of the sublist are at the same state. At this point we use a CAS to set the stCt to -$\infty$ (Line 4). All updates check if the stCt is negative after they increment it. If it is negative, they delegate the request to $S_{2}$. Otherwise, they continue with the operation, delaying the Move until the update increments endCt and so never block client operations.
After stCt is permanently negative (the blue lines other than send replicates reflect this scenario), all updates are redirected towards $S_{2}$, thereby making its sublist version live and leaving the sublist at $S_{1}$ stale. This point begins the Switch operation (or phase). The Switch routine in Algorithm˜5 is invoked to first complete the dynamic re-distribution by updating the next pointer of the previous subtail. Since the current Switch has no control over the background operations on the sublist previous to it, this attempt is done in a loop to account for redirections (expected to not happen more than once, as it is highly undesirable to move the same sublist way too frequently). After this step, the Switch, and hence the re-distribution is completed by updating the registry entry for the sublist in all servers.
258
Move(Entry entry, int newSId):
@@ 5.4 Move, Replay and Switch Algorithms
260  remoteSH = Send MoveSH($\rightarrowsId, head\rightarrowts$, $\rightarrowkeyMax, entry\rightarrow$);
[…]
263  do
264     temp = $\rightarrowendCt + entry\rightarrowoffset$;
265
[…]
267 Switch(entry, newSId);
268
269MoveNext(Ref prev, Ref curr, int newSiD):
@@
275     Send RepDelete($\rightarrownewLoc, curr$, $\rightarrowts, curr\rightarrowsId$) to newSId;
276
@@
278     return;
279
280 MoveItem($\rightarrownewLoc, curr\rightarrownext$, newSId);
281
282Ref MoveSHRecv(int itemSId, int itemTs, Key keyMax):
283  *newstCt = new int[1];
284  *newendCt = new int[1];
285  ST = new Item(ST_KEY, keyMax, , , null, newstCt, newendCt, null);
286  SH = new Item(SH_KEY, $\phi$, itemTs, itemSId, ST, newstCt, newendCt, null);
287  entry = registry.getByKey(keyMax);
@@
292  return SH;
293
[…]
295  while $\rightarrowts \neq prev$Ts $\lor$ $\rightarrowsId \neq prev$SId do
296     $prev\rightarrownext$;
[…]
298        prev = prevNewLoc;
299
300
301 return <<prevNewLoc, Replay(prev, itemTs, key, itemSId, itemTs, false)>>;
302
303Ref RepDeleteRecv(Ref prevNewLoc, int prevSId, int prevTs):
304  Lines 4 to 4;
305  do
[…]
307     ref = temp;
308     ref.mark = true;
309
[…]
311 return prevNewLoc;
312
313Ref MoveItemRecv(Ref prev, Key key, bool isMarked, Ref STNext, int itemSid, int itemTs):
314  if key == ST_KEY
315     curr = prev;
316     do
317        $curr\rightarrownext$;
318
@@
321     return curr;
322
323 return Replay($prev\rightarrowts,$ key, itemSId, itemTs, isMarked);
324
325Ref Replay(Ref prev, Ref compTs, Key key, int itemSId, int itemTs, bool isMarked):
326  do
327     curr = prev;
328     do
329        currPrev = curr;
330        $curr\rightarrownext$;
331
@@
340 return newItem;
341
342InsertReplayResponseRecv(Ref oldLoc, Ref newLoc):
@@
345
346RemoveReplayResponseRecv(Ref oldLoc):
@@
348
Algorithm 4 The Move and Replay Algorithm
349
@@
352     leftEntry = registry.getByKey(keyMin);
@@
360
361       if leftSH == null
362           response = true;
363
364    while $\neg$ response;
@@
366
367 for $i\in serverList-\{me,newSId\}$ do
368     response = Send SwitchServer( keyMax, newSH) to i;
369
370 response = Send SwitchServer(keyMax, newSH) to newSId;
371
372SwitchServerRecv(Key keyMax, Ref newSH:
373  entry = registry.getByKey(keyMax);
@@
375  return true;
376
377Ref SwitchSTRecv(Key keyMin, Ref newSH):
378  leftEntry = registry.getByKey(keyMin);
@@
385
386 return leftSH;
387
388bool switchNextST(Ref leftST, Ref newSH):
@@
395  return true;
396
397
Algorithm 5 Switch Algorithm
6 Correctness
@@
We show the correctness of our algorithms by showing that the client operations are linearizable, that they never get blocked by the background operations. We push the correctness of the background operations to Appendix˜D to save space. However, for the rest of this section, we borrow the following observations proved in the appendix: 1. There is only one active subhead to a sublist at any given time. 2. If an update on a sublist has checked that the sign of stCt is positive after incrementing it, then the stCt of the sublist cannot become negative until the update also increments its endCt.
6.1 Linearizability
@@
We show that the client operations are linearizable by identifying linearization points [19] (LP) for every possible output. DiLi is designed in a way that every operation can be said to take place in some specific machine entirely, by identifying the sublist that the operation takes place in. This is done by noting the currently active subhead for the sublist at the time when the operation is said to be linearized. Thus, our proof sketch for linearizability is by identifying the same linearization points as in the Harris List, while also showing that the same point in time was on the machine with the active sublist.
@@
Find: For a successful Find, the LP is immediately after the execution of the mark check returned false at Line 2 at the machine where Line 2 also had its start count to be non-negative. For an unsuccessful Find, we have two cases – 1. If Line 2 ever was evaluated as true, the LP is after the execution of the key comparison in that line and it took place in the machine where the start count check on the same line returned false. 2. If the above was not satisfied, then a subtail that upper bounds the key must have been reached by the check on Line 2. Since a subtail does not get updated, instead of performing a subsequent start count check here, we instead linearize the find after the execution of Line 2 to be false and the operation took place in the machine where the latter was evaluated to be false.
@@
Remove: For a successful Remove, the LP is immediately after the successful execution of CAS in Line 2, where the next pointer of the item of the matching key gets marked. For an unsuccessful Remove, we have two cases – 1. If the search returned null, The LP is the same as an unsuccessful Find. 2. If the search returned an item of a matching key, the LP is immediately after Line 2 returned true, i.e., when the node is already seen as marked/removed by another concurrent remove of the same key.
@@
Insert: For a successful Insert, the LP is immediately after the successful execution of CAS in Line 3. For an unsuccessful Insert, we similarly have two cases – 1. If the search returned null, the LP is the same as an unsuccessful Find. 2. If the search returned a rightNode of the matching key, the LP is the same as a successful Find.
6.2 Conditional lock-freedom of Client Operations

### v2

@@ 5.2. Client Operations
  end while
end
bool Remove(Key key, Ref SH = null):
  Same as Find but line 1 is changed to
 return Delete(node, key);
  and the corresponding delegates are that of Remove instead of Find.
end
bool Delete(Ref node, Key key):
  result = false;
@@
     return false;
  end if
@@
        break;
     end if
[…]
     newVal = temp;
     newVal.mark = true;
[…]
        result = true;
[…]
       else
[…]
        end if
       break;
     end if
  end while
 leftNode = null;
[…]
  Search(key, SH, leftNode);
  return result;
end
Entry *Split(Entry * entry, Ref sItem):
  *newstCt = new int[1];
  *newendCt = new int[1];
  SH = new Item(SH_KEY, $\phi$, ts.fetch_add(1), me, null, newstCt, newendCt, null);
[…]
 do
[…]
     if temp.mark
        return null;
     end if
@@
  do
     prev = curr;
@@
     $curr\rightarrownext$;
[…]
 do
    $a_{1}$ = newstCt- newendCt ;
     $a_{2}$ = $\rightarrowstCt- sItem\rightarrowendCt$ ;
 while $a_{1}+a_{2}\neq entry\rightarrow offset$ ;
[…]
 registry.addEntry(newEntry);
[…]
 for $i\in(serverList-\{me\})$ do
[…]
  end for
 return newEntry;
end
bool RegisterSublistRecv(Key keyMin, Ref SH):
  Entry leftEntry = registry.getByKey(keyMin);
[…]
  return true;
end
Algorithm 1 Algorithm for $\tt{Find}$, $\tt Insert$}, $\tt{Delete}$} and $\tt Split$} Operations
5.3. The Split Algorithm
The Split operation (Algorithm 1) takes two inputs – a sublist entry and an item (referred to as sItem) at which the sublist is to be split. This truncates the original sublist to a smaller key range and creates a new sublist for the remaining key range. The Split is performed in through 4 major steps, while always guaranteeing that the client operations stay lock-free: (1) Allocate new counters stCt and endCt (lines 1 and 1) for the new sublist. The purpose of these counts is to ensure that if a client operation overlaps with Split/Move, it is detected so that Split/Move can take corrective action. (2) Insert a connected block of a subtail followed by a subhead (as in Lines 1 and 1) at the next pointer of sItem (insertion is in Line 1). If sItem gets deleted before this, split operation fails (Line 1) and can be retried with another call at a different sItem. (3) Update the stCt and endCt pointers of the second half of the sublist (Lines 1 to 1). Note that if an item gets inserted and gets missed in the traversal, it must have copied the counters from the item it was previous to, whose counters (or their ancestors’) were updated during this traversal. Thus, they must also contain the new counter pointers. (4) Update the registry by adding an entry for the second half of the sublist, and then truncating the first half to its reduced keyMax (Lines 1 to 1).
The above approach has two corner cases to handle due to the asynchronous insertion. First, for client operations that concurrently search for a key higher than sItem, the search can get truncated at the inserted subtail in case the client operation traversal encounters it. For this purpose, the subtails contain a keyMax element stored. Specifically, the red lines from 1 to 1 check whether the key belongs to the first half of the sublist that is undergoing a concurrent split or to the second half. If it is in the second half, then its subhead pointer is updated, speeding up potential retries of the lock-free Harris List traversal.
Second, we can no longer expect the difference between stCt and endCt of the individual sublists to be 0 when there are no ongoing updates (insert/remove). This is because, during the counter updates to the second half of the sublist, some updates could have incremented the stCt of the left half and updated the endCt of the right half. This accumulated change to the counter difference when the sublist is static is denoted as the offset of the sublist. As would be explained later, the new offset is important for Move and Switch operations. To compute this new offset, we employ a spin (Lines 1 to 1) that breaks only when the computed values $a_{1}$ and $a_{2}$ reflect the new right half offset and left half offsets respectively. The condition works because the difference between the stCt and endCt of any sublist stabilizes to a minimum only when there are no ongoing updates in the sublist. This is true for both the left and right halves of the sublist, and since the increments went to exactly one of the two stCt counters and one of the two endCt counters for each update, the sum of the new offsets must be equal to the old offset when the two new sublists together have no ongoing updates. We note that this spin does not delay the client operations, and only delays background operations on the split sublists.
5.4. Move, Replay and Switch Algorithms
A Move routine invoked on a server machine (henceforth $S_{1}$) takes the registry entry of a sublist and the machine to move it to (henceforth $S_{2}$), as parameters. The goal of Move is to create a clone of a sublist and then make it the live copy. To achieve this, the Move routine as described in Algorithm 2 traverses each item (by recursively calling the MoveItem routine) in the sublist in $S_{1}$ and creates a copy of them in $S_{2}$. The reference in $S_{2}$ of last moved item is sent back each time so that subsequent items can be linked to the previously moved item.
Move compensates for concurrent updates that the Move routine did not witness during its traversal, by (1) detecting that such concurrent updates occurred, and (2) replicatingthose updates correctly on $S_{2}$. The former is achieved by using stCt (incremented before the start of update operation) and endCt (incremented after the update operation); and the latter is achieved by sending replicates (Line 2 in Algorithm 1 and Line 1 in Algorithm 1) of these operations to $S_{2}$. As discussed in Section 4, the inserts and removes will need further specification for the correct replay of the updates in $S_{2}$, as they may appear out of order when received. To account for this, the replay models the changes to the sublist as a result of the following update operations: InsertAfter(Ref, Item) and Delete(Ref), that distinctly convey where the insert CAS (Line 1) and remove CAS (Line 1) that had succeeded. If multiple items got inserted at a particular Ref $A$, that had been moved by the Move routine, then the item with the lower timestamp got inserted first. As for the Delete, the same item that was moved to $S_{2}$ is the one that gets deleted, and a subsequent insertion for the same key is treated as a different item. Since Ref present in $S_{1}$ for an item is not known by Ref copy of it in $S_{2}$, we denote the uniqueness of an item by the ¡sId, timestamp¿ tuple.
Using the above observations, we developed our Replay algorithm presented in Lines 2 to 2. The received move of an item and a received insert replicate utilize this routine. Once the prev item is identified using their <<sId, timestamp>> tuple, the Replay algorithm takes this prev item and traverses from prev to find the right position to insert the item. To not miss any item, the Move routine even moves marked items as part of the algorithm, which are then delinked once the cloned sublist becomes active. Because replay is concurrently invoked by Move and multiple Replicate threads, insertions at replay are retried. An extensive proof of correctness for the replay is provided in Appendix(Me, 2025b).
Finally, to detect that a sublist is live in $S_{1}$, we utilize the stCt and endCt values of the sublist. All updates increment the stCt before performing the update, and increment the endCt after performing the update. However, if a replicate message is sent, the endCt is incremented only after the completion of the replay (Line 2 and 2). This way, when the endCt measured matches the stCt measured immediately after, we can be sure that both copies of the sublist are at the same state. At this point we use a CAS to set the stCt to -$\infty$ (Line 2). All updates check if the stCt is negative after they increment it. If it is negative, they delegate the request to $S_{2}$. Otherwise, they continue with the operation, delaying the Move until the update increments endCt and so never block client operations.
After stCt is permanently negative (the blue lines other than send replicates reflect this scenario), all updates are redirected towards $S_{2}$, thereby making its sublist version live and leaving the sublist at $S_{1}$ stale. This point begins the Switch operation (or phase). The Switch routine in Algorithm 2 is invoked to first complete the dynamic re-distribution by updating the next pointer of the previous subtail. Since the current Switch has no control over the background operations on the sublist previous to it, this attempt is done in a loop to account for redirections (expected to not happen more than once, as it is highly undesirable to move the same sublist way too frequently). After this step, the Switch, and hence the re-distribution is completed by updating the registry entry for the sublist in all servers.
Move(Entry entry, int newSId):
[…]
  remoteSH = Send MoveSH($\rightarrowsId, head\rightarrowts$, $\rightarrowkeyMax, entry\rightarrow$);
[…]
  do
     temp = $\rightarrowendCt + entry\rightarrowoffset$;
[…]
 Switch(entry, newSId);
end
MoveNext(Ref prev, Ref curr, int newSiD):
[…]
  end if
[…]
     Send RepDelete($\rightarrownewLoc, curr$, $\rightarrowts, curr\rightarrowsId$) to newSId;
  end if
[…]
     return;
  end if
 MoveItem($\rightarrownewLoc, curr\rightarrownext$, newSId);
end
Ref MoveSHRecv(int itemSId, int itemTs, Key keyMax):
  *newstCt = new int[1];
  *newendCt = new int[1];
  ST = new Item(ST_KEY, keyMax, , , null, newstCt, newendCt, null);
  SH = new Item(SH_KEY, $\phi$, itemTs, itemSId, ST, newstCt, newendCt, null);
  entry = registry.getByKey(keyMax);
@@ 5.4. Move, Replay and Switch Algorithms
  while $\rightarrowts \neq prev$Ts $\lor$ $\rightarrowsId \neq prev$SId do
     $prev\rightarrownext$;
[…]
        prev = prevNewLoc;
     end if
  end while
 return <<prevNewLoc, Replay(prev, itemTs, key, itemSId, itemTs, false)>>;
end
Ref RepDeleteRecv(Ref prevNewLoc, int prevSId, int prevTs):
  Lines 2 to 2;
  do
[…]
     ref = temp;
     ref.mark = true;
[…]
 return prevNewLoc;
end
Ref MoveItemRecv(Ref prev, Key key, bool isMarked, Ref STNext, int itemSid, int itemTs):
  if key == ST_KEY
     curr = prev;
     do
        $curr\rightarrownext$;
@@
     return curr;
  end if
 return Replay($prev\rightarrowts,$ key, itemSId, itemTs, isMarked);
end
Ref Replay(Ref prev, Ref compTs, Key key, int itemSId, int itemTs, bool isMarked):
  do
     curr = prev;
     do
        currPrev = curr;
        $curr\rightarrownext$;
@@
[…]
     ref = newItem;
     ref.mark = temp.mark;
[…]
 return newItem;
end
InsertReplayResponseRecv(Ref oldLoc, Ref newLoc):
[…]
end
RemoveReplayResponseRecv(Ref oldLoc):
[…]
end
@@
     leftEntry = registry.getByKey(keyMin);
@@
        end if
       if leftSH == null
           response = true;
        end if
    while $\neg$ response;
  end if
@@
 for $i\in serverList-\{me,newSId\}$ do
     response = Send SwitchServer( keyMax, newSH) to i;
  end for
 response = Send SwitchServer(keyMax, newSH) to newSId;
end
SwitchServerRecv(Key keyMax, Ref newSH:
  entry = registry.getByKey(keyMax);
@@
  return true;
end
Ref SwitchSTRecv(Key keyMin, Ref newSH):
  leftEntry = registry.getByKey(keyMin);
@@
  end if
 return leftSH;
end
bool switchNextST(Ref leftST, Ref newSH):
@@
  return true;
end
Algorithm 2 The Move and Replay Algorithm
6. Correctness
@@
We show the correctness of our algorithms by showing that the client operations are linearizable, and that they never get blocked by the background operations. We push the correctness of the background operations to Appendix(Me, 2025b) to save space. However, for the rest of this section, we borrow the following observations proved in the appendix: (1) There is only one active subhead to a sublist at any given time. (2) If an update on a sublist has checked that the sign of stCt is positive after incrementing it, then the stCt of the sublist cannot become negative until the update also increments its endCt.
6.1. Linearizability
@@
We show that the client operations are linearizable by identifying linearization points (Herlihy and Wing, 1990) (LP) for every possible output. DiLi is designed in a way that every operation can be said to take place in some specific machine entirely, by identifying the sublist that the operation takes place in. This is done by noting the currently active subhead for the sublist at the time when the operation is said to be linearized. Thus, our proof sketch for linearizability is by identifying the same linearization points as in the Harris List, while also showing that the same point in time was on the machine with the active sublist.
@@
Find: For a successful Find, the LP is immediately after the execution of the mark check returned false at Line 1 at the machine where Line 1 also had its start count to be non-negative. For an unsuccessful Find, we have two cases – (1) If Line 1 ever was evaluated as true, the LP is after the execution of the key comparison in that line and it took place in the machine where the start count check on the same line returned false. (2) If the above was not satisfied, then a subtail that upper bounds the key must have been reached by the check on Line 1. Since a subtail does not get updated, instead of performing a subsequent start count check here, we instead linearize the find after the execution of Line 1 to be false and the operation took place in the machine where the latter was evaluated to be false.
@@
Remove: For a successful Remove, the LP is immediately after the successful execution of CAS in Line 1, where the next pointer of the item of the matching key gets marked. For an unsuccessful Remove, we have two cases – (1) If the search returned null, The LP is the same as an unsuccessful Find. (2) If the search returned an item of a matching key, the LP is immediately after Line 1 returned true, i.e., when the node is already seen as marked/removed by another concurrent remove of the same key.
@@
Insert: For a successful Insert, the LP is immediately after the successful execution of CAS in Line 1. For an unsuccessful Insert, we similarly have two cases the LP is the same as a successful Find.
6.2. Conditional lock-freedom of Client Operations